# Weyl semimetallic state with antiferromagnetic order in Rashba-Hubbard model

### v1

@@ III.6 Edge States
As discussed above, there exist WPs inside and on the boundary of reduced Brillouin zone with winding numbers ±$\pm 1$. Therefore, existence of edge states along a quasi-one dimensional chain is guaranteed in a way similar to the case of graphene wakabayashi ; berenvig or other magnetically ordered systems with Dirac points lau ; wang1 . These edge states, in the case of topological insulator, are pairs of states propagating in direction opposite to each other. The edge-state dispersion crosses the Fermi level and connects the valence and conduction bands. In order to explore the edge state in the WSM-AFM state, we consider a ribbon oriented along the $x$-direction consisting of $N$ chains of atoms positioned along $y$ direction such that $k_{x}$ becomes a good quantum number. Thus, the Hamiltonian $H_{Rbx}$ for the ribbon is $4N\times 4N$ matrix given by
@@
In Fig. 9, the bands crossing the Fermi energy and shown in different color are correspond to the edge states. Above and below are the bulk dispersion bands, which are gapped. The edge states cross the Fermi level at four different points, two of the crossings lie very close to $-\pi$ and $\pi$, while the other two lie near $k_{x}=0$. We find that the crossing turns into a flat degenerate bands as the number of chains in the ribbon is increased. Here, we have focussed on the ribbon oriented along $x$ direction. When the ribbon is oriented along the $y$-direction, same edge-state dispersion is obtained, which is the consequence of the reflection symmetry about the line $k_{x}$ = $k_{y}$ in the Brillouin zone. Although we have restricted our calculations to the direction along the primitive translational vectors of the original lattice, it would be interesting to study the edge states if the ribbons are oriented along the primitive translational vector of the reduced Brillouin zone, when the chains consist of sites belonging to a particular sublattice only, i.e., the magnetic moments are aligned along the same direction.
@@ III.7 Quasiparticle interference
The QPI patterns for a non-topological systems is determined primarily by the spectral-density distribution along the constant-energy contours (CECs) for a given energy as well as by the shape of the contours hoffman . In particular, the patterns are dominated by those scattering vectors, which connect the regions with high spectral density roushan . However, the situation is contrastingly different for a helical-Fermi liquid in the topological semimetals, where the backscattering by a non-magnetic impurity atom may not allowed, therefore no characteristic response is expected in the patterns. On the other hand, the magnetic impurity generates only a weak response guo . If the tip of the probe can differentiate between the spin state of the quasiparticle, the QPI patterns can provide crucial information about the band structure in the vicinity of Fermi surface as well as the spin texture. Here, in the WSM-AFM, there exist only two species of quasiparticle, either with spin oriented along $\hat{x}+\hat{y}$ or along -$\hat{x}$ $\hat{y}$, where $\hat{x}$ and $\hat{y}$ are the unit vectors directed along the primitive translational vectors, respectively. The orientation of the quasiparticle spin is in accordance with the direction of the magnetic moments in the WSM-AFM state.
@@
Fig. 10 (a) and (b) show CECs for $\omega=-0.1$ and $-0.2$. There are two types of CECs in the magnetic Brillouin zone. One with the circular shape, which has spins pointing along $\hat{x}+\hat{y}$ in one half and along -$\hat{x}$ $\hat{y}$ in another half. The quasiparticle spin is required to change the direction abruptly at the intersection of a cricular CEC and a line running along the zone-diagonal direction and bisecting the CEC. Another set of CECs appear like the cross section of a banana along its length, with highly non-uniform spectral density along. It may be noted that each of these CECs in the magnetic Brillouin zone has the quasiparticle with spin pointing along only one direction. In other words, upon moving along the CECs and completing one cycle, there is no change in the spin direction. Thus, the orientation of quasiparticle spin changes along CECs, which is different from the way it changes along the cross-section of Dirac cone for in helical liquid existing on the surface of a strong topological insulator guo .
@@
Unlike the Dirac cone generated QPI patterns, where the non-magnetic impurities can give rise to only weak and nonsingular response farrell , the patterns in the WSM-AFM state are not featureless. Fig. 11 shows the QPI patterns obtained for $V_{oi}=0.1$ and $V_{ot}=0.1$. The main reason is only two possible orientation of spins along the CECs. The intrapocket scattering for the banana-shaped CECs generates a pattern of similar shape at $\Gamma$ as the quasiparticle spin has the same orientation all along. The pattern at $\Gamma$ has also intrapocket contribution due to the circular pockets near the points like $(\pi,0)$, which is weak and unnoticeable because of tiny scattering vectors. The patterns at (±$\pm\pi,\mp\pi$) are generated by the interpocket scattering between the banana-shaped CECs belonging to the reduced Brillouin zone and the one outside it, separated by (±$\pm\pi,\mp\pi$) in the momentum space. It may be noted that they have same orientation of spin. On the other hand, the interpocket scattering between the two banana-shaped CECs belonging to the reduced Brillouin zone, does not generate any pattern because of opposite spin orientation. The interpocket scattering between the circular and banana-shaped pocket also creates a noticeable pattern because one side of circular pocket has the quasiparticle spin aligned along the spin of quasiparticle associated with the banana-shaped CECs.
@@
Second row shows $\rho_{01}$, the patterns generated by the magnetic impurity with spin along $x$ direction, when the STM is spin insensitive. The pattern generated by the intrapocket scattering from the banana-shaped CECs appears to be weak. More or less similar behavior is observed for the patterns originating from the interpocket scattering. The weak patterns originate because the quasiparticle spin direction is flipped after impurity scattering. Further, the scattered quasiparticle finds a reduced DOS with spin oriented along positive $x$ direction. Note that this is also true for the quasiparticle having the spin component along the negative $x$ direction. On the other hand, the spin-resolved QPI shows strong patterns, which arise because of the specific orientation of the quasiparticle spin after scattering, and then detected through the spin-sensitive STM probe. $\rho_{11}$ patterns show strong singular features resulting from both the intrapocket and interpocket scatterings. The intrapocket scattering leads to two linear patterns around $\Gamma$ as well as around ±$\pm(\pi,\pm\pi)$. The interpocket scattering between the two nearest banana-shaped CECs results into two linear patterns around (±$\pm\pi/2$, $\mp\pi/2$). The orientation of the pattern follows directly from the alignment of CECs. QPI patterns because of the interpocket scattering between the circular and banana-shaped CECs are also easily noticeable. These features are more or less repeated for $\rho_{12}$ when the impurity spin is oriented along $y$ direction while the STM probe has spin oriented along $x$ direction. The similarity results from the fact that the quasiparticle spins are not oriented along $x$ or $y$ direction instead along $x$ + $y$ or -$x$ $y$.
@@ IV Discussion
In the current work, the phase diagram is obtained at a temperature $T=0.001$ in the unit of $t$. The temperature corresponds to $\sim$ 5 K. A similar phase diagram although without WSM-AFM state has been reported earlier at a very small temperature, when the difference between the free-energy and ground state energy becomes increasingly small kennedy . It be worthwhile to note that when $T\rightarrow 0$ K, the Rashba metallic state will not be found at all. Therefore, the area occupied by the WSM-AFM state will increase upon lowering the temperature. Secondly, with a rise in temperature, the magnetic moments melt within the static mean-field theory, thus, $W_{1}$ will be shifted to the point ($\pi,0$), whereas $W_{2}$ will move toward the momentum with $k_{y_{o}}=\sin^{-1}\big{(}\frac{8t^{2}}{4\lambda^{2}+8t^{2}}\big{)}$. A similar consequence on the phase diagram is expected, when the quantum correction to the sublattice magnetization is incorporated, which will reduce the sublattice magnetization avinash . The current approach ignores the spatial and thermal fluctuations in the magnetic moments. It will be of strong interest to see the consequences of such fluctuations on the stability of WSM state, which can be studied most effectively with the exact-diagonalization + Monte Carlo (ED + MC) approach harun1 ; harun2 .

### v2

@@ III.6 Edge States
As discussed above, there exist WPs inside and on the boundary of reduced Brillouin zone with winding numbers $\pm 1$. Therefore, existence of edge states along a quasi-one dimensional chain is guaranteed in a way similar to the case of graphene wakabayashi ; berenvig or other magnetically ordered systems with Dirac points lau ; wang1 . These edge states, in the case of topological insulator, are pairs of states propagating in direction opposite to each other. The edge-state dispersion crosses the Fermi level and connects the valence and conduction bands. In order to explore the edge state in the WSM-AFM state, we consider a ribbon oriented along the $x$-direction consisting of $N$ chains of atoms positioned along $y$ direction such that $k_{x}$ becomes a good quantum number. Thus, the Hamiltonian $H_{Rbx}$ for the ribbon is $4N\times 4N$ matrix given by
@@
In Fig. 9, the bands crossing the Fermi energy and shown in different color are correspond to the edge states. Above and below are the bulk dispersion bands, which are gapped. The edge states cross the Fermi level at four different points, two of the crossings lie very close to $\pi$ and $\pi$, while the other two lie near $k_{x}=0$. We find that the crossing turns into a flat degenerate bands as the number of chains in the ribbon is increased. Here, we have focused on the ribbon oriented along $x$ direction. When the ribbon is oriented along the $y$-direction, same edge-state dispersion is obtained, which is the consequence of the reflection symmetry about the line $k_{x}$ = $k_{y}$ in the Brillouin zone. Although we have restricted our calculations to the direction along the primitive translational vectors of the original lattice, it would be interesting to study the edge states if the ribbons are oriented along the primitive translational vector of the reduced Brillouin zone, when the chains consist of sites belonging to a particular sublattice only, i.e., the magnetic moments are aligned along the same direction.
@@ III.7 Quasiparticle interference
A conclusive signature of Weyl points may be obtained with the help of spin- and angle-resolved photoemission spectroscopy (ARPES) prx . Besides, the scanning-tunneling microscopy (STM), which can detect the impurity generated modulation in the local density of states, i.e. QPI, can be another effective tool to confirm Weyl points.
@@
The QPI patterns for a non-topological systems is determined primarily by the spectral-density distribution along the constant-energy contours (CECs) for a given energy as well as by the shape of the contours hoffman . In particular, the patterns are dominated by those scattering vectors, which connect the regions with high spectral density roushan . However, the situation is contrastingly different for a helical-Fermi liquid in the topological semimetals, where the backscattering by a non-magnetic impurity atom may not allowed, therefore no characteristic response is expected in the patterns. On the other hand, the magnetic impurity generates only a weak response guo . If the tip of the probe can differentiate between the spin state of the quasiparticle, the QPI patterns can provide crucial information about the band structure in the vicinity of Fermi surface as well as the spin texture. Here, in the WSM-AFM, there exist only two species of quasiparticle, either with spin oriented along $\hat{x}+\hat{y}$ or along -$\hat{x} -$ $\hat{y}$, where $\hat{x}$ and $\hat{y}$ are the unit vectors directed along the primitive translational vectors, respectively. The orientation of the quasiparticle spin is in accordance with the direction of the magnetic moments in the WSM-AFM state.
@@
Fig. 10 (a) and (b) show CECs for $\omega=-0.1$ and 0.2-0.2. There are two types of CECs in the magnetic Brillouin zone. One with the circular shape, which has spins pointing along $\hat{x}+\hat{y}$ in one half and along -$\hat{x} -$ $\hat{y}$ in another half. The quasiparticle spin is required to change the direction abruptly at the intersection of a circular CEC and a line running along the zone-diagonal direction and bisecting the CEC. Another set of CECs appear like the cross section of a banana along its length, with highly non-uniform spectral density along. It may be noted that each of these CECs in the magnetic Brillouin zone has the quasiparticle with spin pointing along only one direction. In other words, upon moving along the CECs and completing one cycle, there is no change in the spin direction. Thus, the orientation of quasiparticle spin changes along CECs, which is different from the way it changes along the cross-section of Dirac cone for in helical liquid existing on the surface of a strong topological insulator guo .
@@
Unlike the Dirac cone generated QPI patterns, where the non-magnetic impurities can give rise to only weak and nonsingular response farrell , the patterns in the WSM-AFM state are not featureless. Fig. 11 shows the QPI patterns obtained for $V_{oi}=0.1$ and $V_{ot}=0.1$. The main reason is only two possible orientation of spins along the CECs. The intrapocket scattering for the banana-shaped CECs generates a pattern of similar shape at $\Gamma$ as the quasiparticle spin has the same orientation all along. The pattern at $\Gamma$ has also intrapocket contribution due to the circular pockets near the points like $(\pi,0)$, which is weak and unnoticeable because of tiny scattering vectors. The patterns at ($\pm\pi,\mp\pi$) are generated by the interpocket scattering between the banana-shaped CECs belonging to the reduced Brillouin zone and the one outside it, separated by ($\pm\pi,\mp\pi$) in the momentum space. It may be noted that they have same orientation of spin. On the other hand, the interpocket scattering between the two banana-shaped CECs belonging to the reduced Brillouin zone, does not generate any pattern because of opposite spin orientation. The interpocket scattering between the circular and banana-shaped pocket also creates a noticeable pattern because one side of circular pocket has the quasiparticle spin aligned along the spin of quasiparticle associated with the banana-shaped CECs.
@@
Second row shows $\rho_{01}$, the patterns generated by the magnetic impurity with spin along $x$ direction, when the STM is spin insensitive. The pattern generated by the intrapocket scattering from the banana-shaped CECs appears to be weak. More or less similar behavior is observed for the patterns originating from the interpocket scattering. The weak patterns originate because the quasiparticle spin direction is flipped after impurity scattering. Further, the scattered quasiparticle finds a reduced DOS with spin oriented along positive $x$ direction. Note that this is also true for the quasiparticle having the spin component along the negative $x$ direction. On the other hand, the spin-resolved QPI shows strong patterns, which arise because of the specific orientation of the quasiparticle spin after scattering, and then detected through the spin-sensitive STM probe. $\rho_{11}$ patterns show strong singular features resulting from both the intrapocket and interpocket scatterings. The intrapocket scattering leads to two linear patterns around $\Gamma$ as well as around $\pm(\pi,\pm\pi)$. The interpocket scattering between the two nearest banana-shaped CECs results into two linear patterns around ($\pm\pi/2$, $\mp\pi/2$). The orientation of the pattern follows directly from the alignment of CECs. QPI patterns because of the interpocket scattering between the circular and banana-shaped CECs are also easily noticeable. These features are more or less repeated for $\rho_{12}$ when the impurity spin is oriented along $y$ direction while the STM probe has spin oriented along $x$ direction. The similarity results from the fact that the quasiparticle spins are not oriented along $x$ or $y$ direction instead along $x$ + $y$ or -$x$ - $y$.
@@ IV Discussion
In the current work, the phase diagram is obtained at a temperature $T=0.001$ in the unit of $t$. This particular value of temperature corresponds to $\sim$ 5 K if $t$ is taken to be $\sim$ 500meV. With rise in the temperature, the magnetic moments can melt away within the static mean-field theory or randomness can be introduced in their orientation if considered within a more sophisticated approaches involving Monte-Carlo simulations based on the auxiliary-field methods harun1 ; harun2 . For this reason, the AFM order makes room for the RM phase in the phase diagram. A similar phase diagram although without WSM-AFM state has been reported earlier at a very small temperature, when the difference between the free-energy and ground state energy becomes increasingly small kennedy . It is worthwhile to note that when $T\rightarrow 0$ K, the Rashba metallic state will not be found at all. Therefore, the area occupied by the WSM-AFM state will increase upon lowering the temperature. Secondly, with a rise in temperature, as the magnetic moments melt away, $W_{1}$ will be shifted to the point ($\pi,0$), whereas $W_{2}$ will move toward the momentum with $k_{y_{o}}=\sin^{-1}\big(\frac{8t^{2}}{4\lambda^{2}+8t^{2}}\big)$. A similar consequence on the phase diagram is expected, when the quantum correction to the sublattice magnetization is incorporated, which will reduce the sublattice magnetization avinash . The current approach ignores the spatial and thermal fluctuations in the magnetic moments. It will be of strong interest to see the consequences of such fluctuations on the stability of WSM state, which can be studied most effectively with the exact-diagonalization + Monte Carlo (ED + MC) approach harun1 ; harun2 .
@@
Here, we have restricted our effort to half filling with a focus on the previously reported metallic antiferromagnetic state with pseudogap-like features in the density of states kennedy . It would be interesting, in future work, to examine the possible existence of topological states with incommensurate magnetic order stabilized away from half filling. Morever, our calculation shows that the second-neigbhor hopping can also support the Weyl points but not necessarily a Weyl semimetallic state. This is because the pairs of the Weyl points are shifted away from the Fermi level in opposite direction. It may appear that the band filling can be changed in order to force one of the Weyl points to lie at the Fermi level. However, it will be also accompanied by additional bands crossing the Fermi level. In addition, the nature of magnetic order may not necessarily remain to be of the checkerboard-type.